OpenStudy (rock_mit182):

Phosphorus trifluoride is a highly toxic gas that reacts slowly with water to give a mixture of phosphorous acid and hydrofluoric acid. (a) Write a balanced chemical equation for this reaction.

OpenStudy (rock_mit182):

Determine the concentration (in moles per liter) of each of the acids that result from the reaction of 1.94 L of phosphorus trifluoride (measured at 25°C and 0.970 atm pressure) with water to give a solution volume of 872 mL.

OpenStudy (anonymous):

A) PF3 + 3H2O - H3PO3 + 3HF B) from my experience, simply base in the reaction's ratio of the moles of all the reactants. This ratio is exactly the same for volume. Indeed, it is important to find the number of moles, using the formula pV=nRT, where R is the gas constant, T is tempertaure in Kelvin, p is pressure in Pascal and V is volume in m^3. Making n (mol) the subject of the formula becomes n=pV/RT. n=97970*0.00194/8.31*298=190.1/2476.4= 0.08 moles. Meaning, if the volume of PF3 is 1.94L and the number of moles is 0.08moles, the volume of H3PO3 is 1.94L and of moles is 0.08mol, because their ratio is 1:1 and that of HF is 1.94*3= 5.82L and 0.08*3= 0.24 moles. Use the formula n=CV, where n is moles, C is concentration and V is volume in Litres or dm^3. Making C the subject, you get C=n/V. For phosphorus acid, it will be C=0.08/1.94 and for HF will be 0.24/5.82=?